magnesium carbonate, MgCO3/ forms a hydrate containing 39.1% water of hydration. Calculate the formula of this hydrate containing 39.1 % water of hydration. Calculate the formula of this hydrate

1
Expert's answer
2022-01-27T12:10:04-0500

M (MgCO3) = 84.3 g/mol

M (H2O) = 18 g/mol

Lets assume, we have 100 g of MgCO3 · xH2O.

39.1 g of this mass refers to water.

100 - 39.1 = 60.9 g of this mass refers to MgCO3.

n (MgCO3) = m / M = 60.9 / 84.3 = 0.72 mol

n (H2O) = 39.1 / 18 = 2.2 mol

MgCO3 : H2O = 0.72 : 2.2 or approximately 1 : 2. Therefore, the formula of this hydrate is MgCO3 · 2H2O.